A Complete Bibliography of Journal of the Royal Statistical Society, Series D (The Statistician):

1950–1959

Nelson H. F. Beebe University of Utah

Department of Mathematics, 110 LCB

155 S 1400 E RM 233 Salt Lake City, UT 84112-0090

USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

E-mail: beebe@math.utah.edu, beebe@acm.org,
beebe@computer.org (Internet)
WWW URL: http://www.math.utah.edu/~beebe/

13 October 2017 Version 1.01

Title word cross-reference

3 [104]. G [282].

-D [104].

14th [40]. 15th [40]. 16th [40]. 1951 [40]. 1956 [242].

Abstract [16, 242]. Accidents [103]. Accountants [241]. Accounting [77, 182, 221]. Accounts [258]. Accurate [127]. Ackoff [154]. Action [120]. Additions [176]. Advertising [59]. Age [6]. Age-Standardized [6]. Agency [69].

Agricultural [277]. Agriculture [43, 209, 161]. Aid [307]. Air [216, 238]. Aitchison [265]. Allen [225, 9]. American [121, 292]. Among [172]. Analysis [39, 72, 309, 249, 281, 310, 186, 61, 149, 297, 297, 194, 121]. Anderson [297]. Annual [84, 136, 217, 7, 27, 64, 57, 106, 90, 240, 227, 237, 284, 259, 274, 318, 242]. Application [281]. Applications [250, 98, 187]. Applied [209, 299]. Approach [107, 225]. Approximation [183]. Area [121]. Arithmetic [51]. Aspects [184, 87, 297, 137, 172]. Assessing [233]. Association [27, 64, 106, 105, 91, 156, 165, 166, 167, 177, 178, 196, 205, 228, 244, 267, 289, 300].

Attitude [282]. Attribute [304]. Audit [108]. Auditors [241].

Back [11, 21, 28, 35, 46, 54, 62, 73, 82, 92, 99, 110, 124, 132, 140, 146, 157, 168, 179, 188, 197, 206, 218, 229, 245, 253, 268, 276, 288, 301, 312, 321]. Background [208]. Backgrounds [120]. Balliol [166, 227]. Bartlett [187]. Bell [70]. Binary [51]. Binomial [183]. Biological [277, 42]. Blackwell [153]. Board [25]. Book [226, 266, 225, 252, 265, 277, 299, 309, 213, 120, 121, 263, 9, 34, 118, 130, 138, 242, 243, 311, 78, 241, 212, 278, 310, 195, 297, 19, 70, 119, 153, 154, 155, 164, 186, 203, 214, 262, 264, 320, 298, 163, 187, 194, 296, 18, 275, 287, 117]. Boyd [243]. Brewing [185]. British [212, 200, 43]. Brown [265]. Buckland [243]. Budgetary [278]. Budgets [186]. Business [60, 209, 201, 66, 114, 15].

C [265, 34, 130, 119, 275]. Calculating [316]. Calculations [115]. Callender [311]. Cambridge [131, 119]. Canteen [202]. Cardpack [19, 163]. Census [96]. Censuses [138, 135, 214, 81]. Central [113, 242]. Chain [234]. Chance [19, 163]. Characteristic [304]. Chemical [152]. Chessboard [19, 163]. Chew [320]. Choice [19, 163]. Circulation [120]. Class [128]. Classification [184, 52]. Clothing [264]. Co [224]. Co-Efficients [224]. Cochran [263]. Coefficients [316]. College [131, 166, 193, 227]. Commercial [87]. Computation [97]. Computers [50, 76]. Conceptional [182]. Conference [40, 131, 166, 193, 64, 106, 84, 240, 227, 284, 259, 14]. Confidence [293]. Connor [226]. Consumer [211]. Contiguity [143]. Control [4, 311]. Corrections [176]. Correlation [224, 95, 104, 316]. Correspondence [10, 20]. Cost [115, 221]. Cotton [151, 144]. Council [264, 33, 45, 53, 58, 75]. Course [287].

Courses [236]. Cox [263]. Crisis [204, 191]. Customers [234].

D [9, 311, 119, 153, 104]. Dairy [249].

Damned [272]. Dangerous [32]. Dangers [129]. Data [256, 184]. David [118]. Davies [252, 309, 278]. December [100]. Decisions [153]. Defence [271]. Delinquency [173, 172]. Demand [149]. Department [69, 234, 17, 25, 5]. Design [30, 154, 208, 222, 235, 273, 76, 195, 309]. Designs [320, 235, 263]. Details [69, 25]. Determining [293]. Development [278, 152]. Developments [311, 292, 98]. Devons [212]. Diagrams [224]. Dictionary [243]. Difficulties [129]. Difficulty [283]. Digital [50]. Distribution [250, 96, 183, 265]. Distributions [256, 223, 43]. Distributive [88]. Documents [30]. Durables [211].

Earners [172]. Economic

[122, 182, 212, 292, 109]. Economists [9]. Economy [200, 171]. Edition [155, 138]. Editorial [3, 13, 22, 29, 37, 47, 55, 63, 74, 83, 93, 101, 111, 125, 133, 141, 147, 158, 169, 180, 189, 198, 207, 219, 230, 246, 254, 269, 279, 290, 302, 313]. Education [215, 239, 251, 285, 295, 308, 319]. Efficiency [233]. Efficients [224]. Electronic [50, 97]. Elegance [32]. Elementary [119]. Elements [130]. End [123]. Engineering [307]. Engineers [299]. Errata [122]. Erratum [204]. Estimation [224, 6, 149]. Europe [81]. Examinations [39, 38, 72, 71, 100, 116, 145, 175, 260]. Excursion [258]. Exercises [164]. Experimental [263, 195, 320]. Experimentation [317, 42]. Experiments [208, 222, 235, 273, 309]. Experts [81]. Exponential [250].

F [277, 34, 118, 310]. Factorial [273]. Facts [120, 120, 120]. Family [186]. Farms [249].

Federer [195]. Fifth [136]. Figureless [258]. Files [120]. Filing [120]. Firm [233]. First [287]. Fisher [277]. Forecasting [60, 209]. Fourth [84]. Frederick [310]. Frequency [43]. Front [1, 12, 26, 36, 44, 49, 56, 67, 79, 86, 94, 102, 112, 126, 134, 142, 148, 160, 170, 181, 190, 199, 210, 220, 231, 247, 255, 270, 280, 291, 303, 314]. Function [88]. Fundamentals [296]. Future [294, 31].

G [263, 9, 34, 130, 243, 52, 155, 164, 298]. Gallup [139]. Games [153]. Gas [25]. General [7, 27, 57, 90, 136, 237, 274, 318]. Generation [223]. Girshick [153]. Goodman [266]. Governments [114]. Graphs [104]. Grassland [249]. Group [81]. Groups [122, 109].

H [226, 186, 262, 296, 18]. Harrison [34]. Harry [225]. Health [162, 176]. Henry [262]. High [76]. Hoel [155]. Hogben [19, 163]. Hollerith [52]. Holmes [117]. Holmstrom [120, 120]. Home [34]. Houthakker [186].

I.P.A. [248]. II [222, 163, 104]. III. [235]. Illustrative [223]. Income [256, 223, 203, 127]. Incorporated [27, 156, 165, 166, 167, 177, 178, 196, 205, 228, 244, 267, 289, 300]. Index [85, 68, 18]. Indexing [120]. Industrial [309, 4, 311, 135, 17, 41, 81, 214]. Industry [185, 129, 144, 320]. Inflation [232]. Information [258]. Inland [256]. Input [249]. Input-Output [249]. Intensity [89]. Intensive [249]. International [114]. Intervals [293]. Introduction [212, 297, 187, 18, 155]. Investigation [275]. Investigations [16]. Investment [80]. IV [273].

J [226, 265, 120, 310, 119, 186, 18]. Japanese [171]. Jesus [193]. John [241]. Johnson [78]. Joint [264]. July [116, 145, 175, 259]. June [39, 38, 72, 71]. Juvenile [173, 172].

Kaj [194]. Kendall [243, 52, 164].

L [226, 252, 309, 241, 154, 194, 117]. Labour [150]. Lambe [130]. Lancelot [19, 163]. Laurence [241]. Lies [272]. Limitations [61]. Limited [166, 27, 165, 177]. Lindley [119]. List [38, 71, 100, 116, 145, 175]. Logical [76]. Lognormal [265]. Loveday [287]. Ltd. [264]. Luncheon [105].

M [263, 52, 153, 164, 187, 296]. Machines [97]. Management [258, 307, 144]. Managerial [23]. Manufacture [174]. Mapping [143]. Market [24, 69, 184, 211, 213, 25, 34]. Marketing [281, 59]. Martin [311]. Materials [306]. Mathematical [155]. Mathematician [95, 104]. Mathematics [70]. Mathematics-Queen [70]. Matter [11, 21, 1, 12, 28, 35, 46, 26, 36, 44, 54, 62, 73, 82, 49, 56, 67, 79, 92, 99, 110, 124, 86, 94, 102, 112, 132, 140, 146, 126, 134, 142, 157, 168, 179, 148, 160, 170, 188, 197, 206, 218, 181, 190, 199, 210, 229, 245, 220, 231, 253, 268, 276, 288, 247, 255, 270, 280, 301, 312, 321, 291, 303, 314]. Maurice [243]. McCarthy [310]. Measurement [89]. Measurements [264, 222]. Measures [107]. Median [293]. Medical [277, 30, 137]. Meeting [7, 27, 57, 90, 136, 217, 237, 274, 318]. Message [2]. Method [293, 317]. Methods [252, 4, 138, 6, 187, 152, 275, 98, 203, 194]. Miller [119]. Ministry [216, 238]. Minutes [217]. Missing [41]. Mitchell [34]. Models [223]. Moore [298]. Morrell [226, 18]. Moser [275]. Motivation [262]. Multivariate [297]. Municipal [48].

N [118, 297]. National [127, 203]. Nature [222]. Neter [241]. Newspaper [315]. Nielson [194]. No [242]. Nomogram [316].

Non [95, 104]. Non-Mathematician [95, 104]. North [25]. Note [293, 89]. Notes [8, 33, 45, 53, 58, 75, 122, 109]. Notice [261]. Notices [286]. Numbers [85, 68, 18]. Numerical [194].

O [252, 309]. Occupations [122, 109]. Office [113, 48, 242]. Official [81]. Oliver [243]. One [120]. Operating [304]. Operation [17]. Opinions [310]. Organization [17, 25, 88, 5, 69]. Other [76]. Our [2]. Output [249]. Oxford [40, 166, 193, 227, 14].

P [119, 155, 298]. Paper [174, 81]. Part [120, 95, 104, 208, 222, 235, 273]. Partial [224, 316]. Pass [38, 71, 100, 145, 175]. Past [116, 232]. Philip [310]. Philips [52]. Poll [139]. Polytechnic [236]. Practical [317, 98]. Practice [226]. Prais [186]. Prehistoric [192]. Present [204, 191]. Presentation [52]. President [2]. Prices [85]. Primer [118]. Principles [298]. Prizes [52]. Problems [248, 139, 161, 149, 76]. Procedure [310]. Process [52, 152]. Processes [187]. Production [252]. Productivity [61, 23]. Professional [117]. Professor [52]. Programming [76]. Pursuit [32].

Qualifications [91]. Quality [283, 4, 306]. Queen [131, 70]. Quenouille [296]. Queues [202].

R [226, 266, 9, 243, 278, 287, 117]. Races [192]. Random [87, 16]. Rapid [224]. Rates [6]. Ratio [143]. Raw [151]. Readership [248, 315]. Readings [213]. Reasoning [296]. Recorded [257]. Reference [187]. Regent [236]. Regional [315]. Relationships [77, 182]. Report [81]. Research [24, 252, 69, 184, 211, 201, 25, 144, 108, 277, 154, 262, 213]. Results [39, 72]. Retail [85, 159, 108]. Revenue [256]. Review [226, 266, 225, 252, 265, 277, 299, 309, 213, 120, 121, 263, 9, 34, 118, 130, 138, 242, 243, 311, 78, 241, 212, 278, 310, 195, 297, 19, 70, 119, 153, 154, 155, 164, 186, 203, 214, 262, 264, 320, 298, 163, 187, 194, 296, 18, 275, 287, 117]. Road [103]. Roberts [225]. Ronald [277]. Roy [297]. Russell [154].

S [297, 186, 187]. Sample [128]. Sampling [115, 87, 304, 16, 241, 305, 52, 6, 161, 138, 310]. Scaling [282]. Schemes [115, 304]. School [123]. Science [70]. Scientific [208]. Second [138, 155]. September [40]. Sequential [115]. Servant [70]. Service [216, 238]. Seventh [217]. Shorter [8]. Simple [95, 235]. Sizes [264]. Social [128, 154, 275]. Society [315]. Socio [122, 109]. Socio-Economic [122, 109]. Some [122, 223, 69, 184, 185, 139, 68, 137, 161, 109, 25, 41, 171, 149, 98, 297]. Sources [120, 203]. Soviet [278]. Special [187]. Speed [76]. Standardized [6]. States [135, 214]. Statistical [204, 238, 77, 113, 69, 48, 281, 107, 4, 118, 242, 243, 143, 191, 297, 119, 153, 298, 296, 17, 152, 25, 97, 88, 5, 98, 172, 252, 277, 241]. Statistician [201, 66, 294]. Statisticians [27, 165, 166, 177, 31, 156, 167, 178, 196, 205, 228, 244, 267, 289, 300]. Statistics [236, 24, 216, 299, 185, 192, 271, 242, 174, 129, 162, 176, 317, 292, 155, 203, 137, 307, 221, 151, 59, 114, 159, 15, 41, 171, 127, 23, 144, 80, 65, 266, 130, 212, 164, 287, 226, 225, 9, 78, 117]. Stephan [310]. Sterling [121]. Stochastic [187]. Store [234]. Strategy [281]. Street [236]. Students [117]. Study [182, 192]. Survey [87, 310, 248, 315, 275]. Surveys [128, 138]. Syllables [257]. System [278]. Systematic [16].

T [195, 297, 70]. Tables [277, 119]. Teach [266]. Teaching [317, 65]. Technical [248]. Technique [282]. Techniques [298].

Tendering [306]. Terms [243]. Tetley [78]. Thames [25]. Their [76]. Theoretical [258, 164]. Theory [226, 204, 191, 153]. Thoughts [68]. Time [257]. Tools [4]. Trade [159]. Turnover [150]. Two [120, 95].

Uncertain [306]. United [214, 135]. Use [59, 23]. Useful [127]. Uses [185].

V [225, 119, 320]. Vance [241]. Variables [305, 95]. Volk [299]. Volume [19, 163].

W [225, 263, 278, 195, 297]. Wage [172]. Wage-Earners [172]. Wages [232]. Wallis [225]. Week [123]. Week-End [123]. Weights [107]. Western [81]. William [299, 243]. Women [264]. Work [113, 48]. Workshops [307].

Yates [277]. Yourself [266].

References

Anonymous:1950:FMa

[1] Anonymous. Front matter. The Incorporated Statistician, 1(1):1–28, July 1950. CODEN ???? ISSN 1466-9404.

URL http://www.jstor.org/stable/
2988336.

Beveridge:1950:MOP

[2] Lord Beveridge. A message from our President. The Incorporated Statistician, 1(1):3, July 1950. CODEN ???? ISSN 1466-9404. URL http://www.

jstor.org/stable/2988337.

Anonymous:1950:Ea

[3] Anonymous. Editorial. The Incorporated Statistician, 1(1):4, July 1950. CODEN ???? ISSN 1466-9404. URL http://

www.jstor.org/stable/2988338.

Dudding:1950:SMI

[4] B. P. Dudding. Statistical methods as industrial tools “quality control”. The Incorporated Statistician, 1(1):5–9, July 1950. CODEN ???? ISSN 1466-9404.

URL http://www.jstor.org/stable/
2988339.

Wharton:1950:OSD

[5] A. S. Wharton. The organization of a statistical department. The Incorporated Statistician, 1(1):10–16, July 1950. CO-DEN ???? ISSN 1466-9404. URL http:

//www.jstor.org/stable/2988340.

Morrell:1950:EAS

[6] A. J. H. Morrell. The estimation of agestandardized rates by sampling methods. The Incorporated Statistician, 1(1): 17–20, July 1950. CODEN ???? ISSN

1466-9404. URL http://www.jstor.
org/stable/2988341.

Anonymous:1950:AGM

[7] Anonymous. Annual General Meeting. The Incorporated Statistician, 1(1):21 + 23, July 1950. CODEN ???? ISSN

1466-9404. URL http://www.jstor.
org/stable/2988342.

Anonymous:1950:SN

[8] Anonymous. Shorter notes. The Incorporated Statistician, 1(1):25–26, July 1950. CODEN ???? ISSN 1466-9404.

URL http://www.jstor.org/stable/
2988343.

G:1950:BRB

[9] C. O. G. Book review: Statistics for Economists, by R. G. D. Allen. The Incorporated Statistician, 1(1):27, July 1950. CODEN ???? ISSN 1466-9404.

URL http://www.jstor.org/stable/
2988344.

Morton:1950:C

[10] J. E. Morton, G. B. Oakland, J. R. H. Shaul, and W. J. Heard. Correspondence. The Incorporated Statistician, 1(1):29–31, July 1950. CODEN ???? ISSN 1466-9404. URL http://www.

jstor.org/stable/2988345.

Anonymous:1950:BMa

[11] Anonymous. Back matter. The Incorporated Statistician, 1(1):??, July 1950. CODEN ???? ISSN 1466-9404.

URL http://www.jstor.org/stable/
2988346.

Anonymous:1950:FMb

[12] Anonymous. Front matter. The Incorporated Statistician, 1(2):1–28, December 1950. CODEN ???? ISSN 1466-9404.

URL http://www.jstor.org/stable/
2986623.

Anonymous:1950:Eb

[13] Anonymous. Editorial. The Incorporated Statistician, 1(2):3, December 1950. CODEN ???? ISSN 1466-9404.

URL http://www.jstor.org/stable/
2986624.

S:1950:OC

[14] B. B. S. The Oxford Conference. The Incorporated Statistician, 1(2):4–6, December 1950. CODEN ???? ISSN

1466-9404. URL http://www.jstor.
org/stable/2986625.

Ryan:1950:BS

[15] John Ryan. Business statistics. The Incorporated Statistician, 1(2):7–14, December 1950. CODEN ???? ISSN

1466-9404. URL http://www.jstor.
org/stable/2986626.

Finney:1950:AIR

[16] D. J. Finney. Abstract:—investigations in random and systematic sampling. The Incorporated Statistician, 1(2):15– 20, December 1950. CODEN ???? ISSN

1466-9404. URL http://www.jstor.
org/stable/2986627.

Read:1950:OOI

[17] D. R. Read. The organization and operation of an industrial statistical department. The Incorporated Statistician, 1(2):20–27, December 1950. CODEN ???? ISSN 1466-9404. URL http://

www.jstor.org/stable/2986628.

S:1950:BRB

[18] B. B. S. Book review: An Introduction to Index Numbers, by A. J. H. Morrell. The Incorporated Statistician, 1(2): 29, December 1950. CODEN ???? ISSN

1466-9404. URL http://www.jstor.
org/stable/2986629.

M:1950:BRB

[19] A. J. H. M. Book review: Chance and Choice, by Cardpack and Chessboard. Volume I by Lancelot Hogben. The Incorporated Statistician, 1(2):29–30, December 1950. CODEN ???? ISSN

1466-9404. URL http://www.jstor.
org/stable/2986630.

Turner:1950:C

[20] N. C. Turner. Correspondence. The Incorporated Statistician, 1(2):30, December 1950. CODEN ???? ISSN 1466-

9404. URL http://www.jstor.org/
stable/2986631.

Anonymous:1950:BMb

[21] Anonymous. Back matter. The Incorporated Statistician, 1(2):31, December 1950. CODEN ???? ISSN 1466-9404.

URL http://www.jstor.org/stable/
2986632.

Anonymous:1951:Ea

[22] Anonymous. Editorial. The Incorporated Statistician, 2(1):1, April 1951. CODEN ???? ISSN 1466-9404. URL http://

www.jstor.org/stable/2986388.

Tippett:1951:PSM

[23] L. H. C. Tippett. Productivity statistics for managerial use. The Incorporated Statistician, 2(1):2–9, April 1951. CO-DEN ???? ISSN 1466-9404. URL http:

//www.jstor.org/stable/2986389.

Abrams:1951:SMR

[24] Mark Abrams. Statistics for market research. The Incorporated Statistician, 2(1):10–16, April 1951. CODEN ???? ISSN 1466-9404. URL http://www.

jstor.org/stable/2986390.

Reece:1951:OSD

[25] J. A. Reece. Organization of a statistical department: Some details of market research by The North Thames Gas Board. The Incorporated Statistician, 2(1):17–28, April 1951. CODEN ???? ISSN 1466-9404. URL http://www.

jstor.org/stable/2986391.

Anonymous:1951:FMa

[26] Anonymous. Front matter. The Incorporated Statistician, 2(1):29, April 1951. CODEN ???? ISSN 1466-9404.

URL http://www.jstor.org/stable/
2986387.

Anonymous:1951:AIS

[27] Anonymous. Association of Incorporated Statisticians Limited: Annual General Meeting. The Incorporated Statistician, 2(1):30, April 1951. CO-DEN ???? ISSN 1466-9404. URL http:

                                        //www.jstor.org/stable/2986392.

Anonymous:1951:BMa

[28] Anonymous. Back matter. The Incorporated Statistician, 2(1):??, April 1951. CODEN ???? ISSN 1466-9404.

                                        URL http://www.jstor.org/stable/
                                        2986393.

Anonymous:1951:Eb

[29] Anonymous. Editorial. The Incorporated Statistician, 2(2):1, August 1951. CO-DEN ???? ISSN 1466-9404. URL http:

                                        //www.jstor.org/stable/2986596.

Hogben:1951:DMD

[30] Lancelot Hogben. Design of medical documents. The Incorporated Statistician, 2(2):2–14, August 1951. CODEN ???? ISSN 1466-9404. URL http://

                                        www.jstor.org/stable/2986597.

George:1951:FS

[31] C. Oswald George. The future of statisticians. The Incorporated Statistician, 2 (2):15–22, August 1951. CODEN ???? ISSN 1466-9404. URL http://www.

                                        jstor.org/stable/2986598.

Williams:1951:DPE

[32] G. Prys Williams. On the dangerous pursuit of elegance. The Incorporated Statistician, 2(2):23–30, August 1951. CODEN ???? ISSN 1466-9404.

                                        URL http://www.jstor.org/stable/
                                        2986599.

Anonymous:1951:CNa

[33] Anonymous. Council notes. The Incorporated Statistician, 2(2):31, August 1951. CODEN ???? ISSN 1466-9404.

URL http://www.jstor.org/stable/
2986600.

G:1951:BRB

[34] C. O. G. Book review: The Home Market, by G. Harrison; F. C. Mitchell. The Incorporated Statistician, 2(2):32, August 1951. CODEN ???? ISSN 1466-

9404. URL http://www.jstor.org/
stable/2986601.

Anonymous:1951:BMb

[35] Anonymous. Back matter. The Incorporated Statistician, 2(2):??, August 1951. CODEN ???? ISSN 1466-9404.

URL http://www.jstor.org/stable/
2986602.

Anonymous:1951:FMb

[36] Anonymous. Front matter. The Incorporated Statistician, 2(2):??, August 1951. CODEN ???? ISSN 1466-9404.

URL http://www.jstor.org/stable/
2986595.

Anonymous:1951:Ec

[37] Anonymous. Editorial. The Incorporated Statistician, 2(3):1, November 1951. CODEN ???? ISSN 1466-9404.

URL http://www.jstor.org/stable/
2986634.

Anonymous:1951:JEP

[38] Anonymous. June, 1951 examinations pass list. The Incorporated Statistician, 2(3):2, November 1951. CODEN ???? ISSN 1466-9404. URL http://www.

jstor.org/stable/2986635.

Anonymous:1951:JEA

[39] Anonymous. June, 1951 examinations analysis of results. The Incorporated Statistician, 2(3):3, November 1951. CODEN ???? ISSN 1466-9404.

                                        URL http://www.jstor.org/stable/
                                        2986636.

Anonymous:1951:OCS

[40] Anonymous. The Oxford Conference, 14th, 15th, and 16th September, 1951. The Incorporated Statistician, 2(3):4, November 1951. CODEN ???? ISSN

                                        1466-9404. URL http://www.jstor.
                                        org/stable/2986637.

Ryan:1951:SMI

[41] John Ryan. Some missing industrial statistics. The Incorporated Statistician, 2(3):5–20, November 1951. CODEN ???? ISSN 1466-9404. URL http://

                                        www.jstor.org/stable/2986638.

Wishart:1951:BE

[42] John Wishart. Biological experimentation. The Incorporated Statistician, 2(3):21–29, November 1951. CODEN ???? ISSN 1466-9404. URL http://

                                        www.jstor.org/stable/2986639.

Britton:1951:FDB

[43] D. K. Britton. Frequency distributions in British agriculture. The Incorporated Statistician, 2(3):30–44, November 1951. CODEN ???? ISSN 1466-9404.

                                        URL http://www.jstor.org/stable/
                                        2986640.

Anonymous:1951:FMc

[44] Anonymous. Front matter. The Incorporated Statistician, 2(3):45, November 1951. CODEN ???? ISSN 1466-9404.

URL http://www.jstor.org/stable/
2986633.

Anonymous:1951:CNb

[45] Anonymous. Council notes. The Incorporated Statistician, 2(3):46–48, November 1951. CODEN ???? ISSN 1466-

9404. URL http://www.jstor.org/
stable/2986641.

Anonymous:1951:BMc

[46] Anonymous. Back matter. The Incorporated Statistician, 2(3):??, November 1951. CODEN ???? ISSN 1466-9404.

URL http://www.jstor.org/stable/
2986642.

Anonymous:1952:Ea

[47] Anonymous. Editorial. The Incorporated Statistician, 3(1):1–2, February 1952. CODEN ???? ISSN 1466-9404.

URL http://www.jstor.org/stable/
2986588.

Charles:1952:WMS

[48] Enid Charles. The work of a municipal statistical Office. The Incorporated Statistician, 3(1):3–18, February 1952. CODEN ???? ISSN 1466-9404.

URL http://www.jstor.org/stable/
2986589.

Anonymous:1952:FMa

[49] Anonymous. Front matter. The Incorporated Statistician, 3(1):19–31, February 1952. CODEN ???? ISSN 1466-9404.

URL http://www.jstor.org/stable/
2986587.

Bowden:1952:EDC

[50] Dr. Bowden. Electronic digital computers. The Incorporated Statistician, 3(1): 20–30 + 32–34, February 1952. CODEN ???? ISSN 1466-9404. URL http://

                                        www.jstor.org/stable/2986590.

Michaelson:1952:BA

[51] R. L. Michaelson. Binary arithmetic. The Incorporated Statistician, 3(1):35– 40, February 1952. CODEN ???? ISSN

                                        1466-9404. URL http://www.jstor.
                                        org/stable/2986591.

Kendall:1952:PPH

[52] M. G. Kendall. Presentation of Philips and Hollerith prizes: Professor M. G. Kendall on “A Classification of the Sampling Process”. The Incorporated Statistician, 3(1):41–47, February 1952. CO-DEN ???? ISSN 1466-9404. URL http:

                                        //www.jstor.org/stable/2986592.

Anonymous:1952:CNa

[53] Anonymous. Council notes. The Incorporated Statistician, 3(1):48, February 1952. CODEN ???? ISSN 1466-9404.

                                        URL http://www.jstor.org/stable/
                                        2986593.

Anonymous:1952:BMa

[54] Anonymous. Back matter. The Incorporated Statistician, 3(1):??, February 1952. CODEN ???? ISSN 1466-9404.

                                        URL http://www.jstor.org/stable/
                                        2986594.

Anonymous:1952:Eb

[55] Anonymous. Editorial. The Incorporated Statistician, 3(2):1, June 1952. CODEN ???? ISSN 1466-9404. URL http://

                                        www.jstor.org/stable/2986604.

Anonymous:1952:FMb

[56] Anonymous. Front matter. The Incorporated Statistician, 3(2):2–38, June 1952. CODEN ???? ISSN 1466-9404.

URL http://www.jstor.org/stable/
2986603.

Anonymous:1952:AGM

[57] Anonymous. Annual General Meeting. The Incorporated Statistician, 3(2):3–5, June 1952. CODEN ???? ISSN 1466-

9404. URL http://www.jstor.org/
stable/2986605.

Anonymous:1952:CNb

[58] Anonymous. Council notes. The Incorporated Statistician, 3(2):5, June 1952. CODEN ???? ISSN 1466-9404.

URL http://www.jstor.org/stable/
2986606.

Rodgers:1952:USA

[59] John Rodgers. The use of statistics in advertising and marketing. The Incorporated Statistician, 3(2):6–22, June 1952. CODEN ???? ISSN 1466-9404.

URL http://www.jstor.org/stable/
2986607.

Friday:1952:BF

[60] Frank A. Friday. Business forecasting. The Incorporated Statistician, 3(2):25– 37, June 1952. CODEN ???? ISSN

1466-9404. URL http://www.jstor.
org/stable/2986608.

Smith:1952:LPA

[61] A. R. Smith. The limitations of productivity analysis. The Incorporated Statistician, 3(2):39–56, June 1952. CODEN ???? ISSN 1466-9404. URL http://

www.jstor.org/stable/2986609.

Anonymous:1952:BMb

[62] Anonymous. Back matter. The Incorporated Statistician, 3(2):??, June 1952. CODEN ???? ISSN 1466-9404.

                                        URL http://www.jstor.org/stable/
                                        2986610.

Anonymous:1952:Ec

[63] Anonymous. Editorial. The Incorporated Statistician, 3(3):1, October 1952. CO-DEN ???? ISSN 1466-9404. URL http:

                                        //www.jstor.org/stable/2986395.

Anonymous:1952:ACA

[64] Anonymous. Annual Conference of the Association. The Incorporated Statistician, 3(3):2, October 1952. CODEN ???? ISSN 1466-9404. URL http://

                                        www.jstor.org/stable/2986396.

Wishart:1952:TS

[65] John Wishart. The teaching of statistics. The Incorporated Statistician, 3(3): 3–10, October 1952. CODEN ???? ISSN

                                        1466-9404. URL http://www.jstor.
                                        org/stable/2986397.

Lloyd:1952:SB

[66] Edward L. Lloyd. The statistician in business. The Incorporated Statistician, 3(3):11–23, October 1952. CODEN ???? ISSN 1466-9404. URL http://www.

                                        jstor.org/stable/2986398.

Anonymous:1952:FMc

[67] Anonymous. Front matter. The Incorporated Statistician, 3(3):24–36, October 1952. CODEN ???? ISSN 1466-9404.

                                        URL http://www.jstor.org/stable/
                                        2986394.

Morrell:1952:STI

[68] A. J. H. Morrell. Some thoughts on index numbers. The Incorporated Statistician, 3(3):25–34, October 1952. CODEN ???? ISSN 1466-9404. URL http://

                                        www.jstor.org/stable/2986399.

Cauter:1952:OSD

[69] Tom Cauter. “Organization of a statistical department”: Some details of a market research agency. The Incorporated Statistician, 3(3):37–54, October 1952. CODEN ???? ISSN 1466-9404.

URL http://www.jstor.org/stable/
2986400.

M:1952:BRB

[70] A. J. H. M. Book review: Mathematics-Queen and Servant of Science, by E. T. Bell. The Incorporated Statistician, 3(3): 54, October 1952. CODEN ???? ISSN

1466-9404. URL http://www.jstor.
org/stable/2986401.

Anonymous:1952:JEP

[71] Anonymous. June, 1952 examinations pass list. The Incorporated Statistician, 3(3):55, October 1952. CODEN ???? ISSN 1466-9404. URL http://www.

jstor.org/stable/2986402.

Anonymous:1952:JEA

[72] Anonymous. June, 1952 examinations analysis of results. The Incorporated Statistician, 3(3):56, October 1952. CO-DEN ???? ISSN 1466-9404. URL http:

//www.jstor.org/stable/2986403.

Anonymous:1952:BMc

[73] Anonymous. Back matter. The Incorporated Statistician, 3(3):??, October 1952. CODEN ???? ISSN 1466-9404.

URL http://www.jstor.org/stable/
2986404.

Anonymous:1952:Ed

[74] Anonymous. Editorial. The Incorporated Statistician, 3(4):1, December 1952. CODEN ???? ISSN 1466-9404.

                                        URL http://www.jstor.org/stable/
                                        2986504.

Anonymous:1952:CNc

[75] Anonymous. Council notes. The Incorporated Statistician, 3(4):2, December 1952. CODEN ???? ISSN 1466-

                                        9404. URL http://www.jstor.org/
                                        stable/2986505.

Womersley:1952:HSC

[76] J. R. Womersley. High speed computers: Their logical design, programming and other problems. The Incorporated Statistician, 3(4):3–14, December 1952. CO-DEN ???? ISSN 1466-9404. URL http:

                                        //www.jstor.org/stable/2986506.

Bray:1952:ASR

[77] F. Sewell Bray. Accounting and statistical relationships. The Incorporated Statistician, 3(4):15–35, December 1952. CODEN ???? ISSN 1466-9404.

                                        URL http://www.jstor.org/stable/
                                        2986507.

H:1952:BRB

[78] J. H. Book review: Statistics, by Johnson; Tetley. The Incorporated Statistician, 3(4):35, December 1952. CODEN ???? ISSN 1466-9404. URL http://

                                        www.jstor.org/stable/2986508.

Anonymous:1952:FMd

[79] Anonymous. Front matter. The Incorporated Statistician, 3(4):36–54, December 1952. CODEN ???? ISSN 1466-9404.

                                        URL http://www.jstor.org/stable/
                                        2986503.

Whyte:1952:SI

[80] Lewis G. Whyte. Statistics and investment. The Incorporated Statistician, 3(4):37–52, December 1952. CODEN ???? ISSN 1466-9404. URL http://

www.jstor.org/stable/2986509.

S:1952:OPI

[81] B. B. S. Official paper: Industrial censuses in Western Europe, report of a group of experts. The Incorporated Statistician, 3(4):55–56, December 1952. CODEN ???? ISSN 1466-9404.

URL http://www.jstor.org/stable/
2986510.

Anonymous:1952:BMd

[82] Anonymous. Back matter. The Incorporated Statistician, 3(4):??, December 1952. CODEN ???? ISSN 1466-9404.

URL http://www.jstor.org/stable/
2986511.

Anonymous:1953:Ea

[83] Anonymous. Editorial. The Incorporated Statistician, 4(1):1, March 1953. CO-DEN ???? ISSN 1466-9404. URL http:

//www.jstor.org/stable/2986513.

Anonymous:1953:FAC

[84] Anonymous. Fourth Annual Conference. The Incorporated Statistician, 4 (1):2, March 1953. CODEN ???? ISSN

1466-9404. URL http://www.jstor.
org/stable/2986514.

Allen:1953:INR

[85] R. G. D. Allen. Index numbers of retail prices. The Incorporated Statistician, 4(1):3–6, March 1953. CODEN ???? ISSN 1466-9404. URL http://

www.jstor.org/stable/2986515.

Anonymous:1953:FMa

[86] Anonymous. Front matter. The Incorporated Statistician, 4(1):7–41, March 1953. CODEN ???? ISSN 1466-9404.

URL http://www.jstor.org/stable/
2986512.

Edwards:1953:ARS

[87] Frederick Edwards. Aspects of random sampling for a commercial survey. The Incorporated Statistician, 4(1):9–26, March 1953. CODEN ???? ISSN

1466-9404. URL http://www.jstor.
org/stable/2986516.

Turner:1953:SFD

[88] N. C. Turner. The statistical function in a distributive organization. The Incorporated Statistician, 4(1):27–40, March 1953. CODEN ???? ISSN 1466-9404.

URL http://www.jstor.org/stable/
2986517.

Harris:1953:NIM

[89] Amelia Harris and Leslie T. Wilkins. A note on intensity measurement. The Incorporated Statistician, 4(1):42–49, March 1953. CODEN ???? ISSN

1466-9404. URL http://www.jstor.
org/stable/2986518.

Anonymous:1953:AGM

[90] Anonymous. Annual General Meeting. The Incorporated Statistician, 4(1):50– 51, March 1953. CODEN ???? ISSN

1466-9404. URL http://www.jstor.
org/stable/2986519.

Anonymous:1953:AQ

[91] Anonymous. Association qualifications. The Incorporated Statistician, 4(1):52, March 1953. CODEN ???? ISSN

1466-9404. URL http://www.jstor.
org/stable/2986520.

Anonymous:1953:BMa

[92] Anonymous. Back matter. The Incorporated Statistician, 4(1):??, March 1953. CODEN ???? ISSN 1466-9404.

URL http://www.jstor.org/stable/
2986521.

Anonymous:1953:Eb

[93] Anonymous. Editorial. The Incorporated Statistician, 4(2):53, August 1953. CO-DEN ???? ISSN 1466-9404. URL http:

//www.jstor.org/stable/2986443.

Anonymous:1953:FMb

[94] Anonymous. Front matter. The Incorporated Statistician, 4(2):54–61, August 1953. CODEN ???? ISSN 1466-9404.

URL http://www.jstor.org/stable/
2986442.

Keen:1953:CNMa

[95] Joan Keen and D. J. Page. Correlation for the non-mathematician. Part I. simple correlation of two variables. The Incorporated Statistician, 4(2):55–60 + 62– 65, August 1953. CODEN ???? ISSN

1466-9404. URL http://www.jstor.
org/stable/2986444.

Mason:1953:CD

[96] J. I. Mason. The census of distribution. The Incorporated Statistician, 4(2):66– 80, August 1953. CODEN ???? ISSN

1466-9404. URL http://www.jstor.
org/stable/2986445.

Swann:1953:SCE

[97] B. B. Swann. Statistical computation and electronic machines. The Incorporated Statistician, 4(2):81–91, August 1953. CODEN ???? ISSN 1466-9404.

URL http://www.jstor.org/stable/
2986446.

Wilkins:1953:PAS

[98] Leslie T. Wilkins. Practical applications of some recent developments in statistical methods. The Incorporated Statistician, 4(2):92–106, August 1953. CO-DEN ???? ISSN 1466-9404. URL http:

                                        //www.jstor.org/stable/2986447.

Anonymous:1953:BMb

[99] Anonymous. Back matter. The Incorporated Statistician, 4(2):107, August 1953. CODEN ???? ISSN 1466-9404.

                                        URL http://www.jstor.org/stable/
                                        2986449.

Anonymous:1953:DEP

[100] Anonymous. December, 1952, examinations pass list. The Incorporated Statistician, 4(2):108, August 1953. CODEN ???? ISSN 1466-9404. URL http://

                                        www.jstor.org/stable/2986448.

Anonymous:1953:Ec

[101] Anonymous. Editorial. The Incorporated Statistician, 4(3):109, October 1953. CODEN ???? ISSN 1466-9404.

                                        URL http://www.jstor.org/stable/
                                        2988327.

Anonymous:1953:FMc

[102] Anonymous. Front matter. The Incorporated Statistician, 4(3):110–147, October 1953. CODEN ???? ISSN 1466-

                                        9404. URL http://www.jstor.org/
                                        stable/2988326.

Williams:1953:RA

[103] G. Prys Williams. Road accidents. The Incorporated Statistician, 4(3):111–127, October 1953. CODEN ???? ISSN

                                        1466-9404. URL http://www.jstor.
                                        org/stable/2988328.

Keen:1953:CNMb

[104] Joan Keen and A. N. James. Correlation for the non-mathematician. Part II. 3-D graphs. The Incorporated Statistician, 4(3):128–130 + 134–135, October 1953. CODEN ???? ISSN 1466-9404.

URL http://www.jstor.org/stable/
2988329.

Anonymous:1953:AL

[105] Anonymous. Association luncheon, 1953. The Incorporated Statistician, 4 (3):131–133, October 1953. CODEN ???? ISSN 1466-9404. URL http://

www.jstor.org/stable/2988330.

Anonymous:1953:ACA

[106] Anonymous. The Annual Conference of the Association, 1953. The Incorporated Statistician, 4(3):136–137, October 1953. CODEN ???? ISSN 1466-

9404. URL http://www.jstor.org/
stable/2988331.

Desmond:1953:SAW

[107] D. J. Desmond. The statistical approach to weights and measures. The Incorporated Statistician, 4(3):138–146, October 1953. CODEN ???? ISSN 1466-

9404. URL http://www.jstor.org/
stable/2988332.

Treasure:1953:RAR

[108] J. A. P. Treasure. Retail audit research.

The Incorporated Statistician, 4(3):148– 159, October 1953. CODEN ???? ISSN

1466-9404. URL http://www.jstor.
org/stable/2988333.

Parkinson:1953:SNO

[109] Lindsay Parkinson. Some notes on occupations and socio-economic groups. The Incorporated Statistician, 4(3):160–163, October 1953. CODEN ???? ISSN

                                        1466-9404. URL http://www.jstor.
                                        org/stable/2988334. See errata [122].

Anonymous:1953:BMc

[110] Anonymous. Back matter. The Incorporated Statistician, 4(3):164, October 1953. CODEN ???? ISSN 1466-9404.

                                        URL http://www.jstor.org/stable/
                                        2988335.

Anonymous:1953:Ed

[111] Anonymous. Editorial. The Incorporated Statistician, 4(4):165, December 1953. CODEN ???? ISSN 1466-9404.

                                        URL http://www.jstor.org/stable/
                                        2986523.

Anonymous:1953:FMd

[112] Anonymous. Front matter. The Incorporated Statistician, 4(4):166–208, December 1953. CODEN ???? ISSN

                                        1466-9404. URL http://www.jstor.
                                        org/stable/2986522.

Campion:1953:WCS

[113] H. Campion. The work of the central statistical Office. The Incorporated Statistician, 4(4):167–177, December 1953. CODEN ???? ISSN 1466-

                                        9404. URL http://www.jstor.org/
                                        stable/2986524.

Rodgers:1953:ISG

[114] John Rodgers and J. I. Mason. International statistics for governments and business. The Incorporated Statistician, 4(4):179–195, December 1953. CODEN ???? ISSN 1466-9404. URL http://

                                        www.jstor.org/stable/2986525.

Champernowne:1953:CCS

[115] D. G. Champernowne. Cost calculations for sequential sampling schemes. The Incorporated Statistician, 4(4):197–207, December 1953. CODEN ???? ISSN

1466-9404. URL http://www.jstor.
org/stable/2986526.

Anonymous:1953:JEP

[116] Anonymous. July, 1953 examinations past list. The Incorporated Statistician, 4(4):209–211, December 1953. CODEN ???? ISSN 1466-9404. URL http://

www.jstor.org/stable/2986527.

W:1953:BRB

[117] L. T. W. Book review: Statistics for Professional Students, by R. L. A. Holmes. The Incorporated Statistician, 4(4):212, December 1953. CODEN ???? ISSN 1466-9404. URL http://www.

jstor.org/stable/2986528.

G:1953:BRB

[118] H. W. G. G. Book review: A Statistical Primer, by F. N. David. The Incorporated Statistician, 4(4):212–213, December 1953. CODEN ???? ISSN

1466-9404. URL http://www.jstor.
org/stable/2986529.

M:1953:BRB

[119] A. J. H. M. Book review: Cambridge Elementary Statistical Tables, by D. V. Lindley; J. C. P. Miller. The Incorporated Statistician, 4(4):213–214, December 1953. CODEN ???? ISSN

1466-9404. URL http://www.jstor.
org/stable/2986530.

F:1953:BRBa

[120] F. A. F. Book review: Facts, Files and Action. Part One: Sources and Backgrounds of Facts, by J. E. Holmstrom; Facts, Files and Action. Part Two: Filing, Indexing, and Circulation, by J. E.

Holmstrom. The Incorporated Statistician, 4(4):214, December 1953. CODEN ???? ISSN 1466-9404. URL http://

                                        www.jstor.org/stable/2986531.

F:1953:BRBb

[121] F. A. F. Book review: The Sterling Area: An American Analysis. The Incorporated Statistician, 4(4):215, December 1953. CODEN ???? ISSN 1466-9404.

                                        URL http://www.jstor.org/stable/
                                        2986532.

Anonymous:1953:ESN

[122] Anonymous. Errata: Some notes on occupations and socio-economic groups. The Incorporated Statistician, 4(4):216, December 1953. CODEN ???? ISSN

                                        1466-9404. URL http://www.jstor.
                                        org/stable/2986534. See [109].

Anonymous:1953:WES

[123] Anonymous. Week-end school. The Incorporated Statistician, 4(4):216, December 1953. CODEN ???? ISSN

                                        1466-9404. URL http://www.jstor.
                                        org/stable/2986533.

Anonymous:1953:BMd

[124] Anonymous. Back matter. The Incorporated Statistician, 4(4):??, December 1953. CODEN ???? ISSN 1466-9404.

                                        URL http://www.jstor.org/stable/
                                        2986535.

Anonymous:1954:Ea

[125] Anonymous. Editorial. The Incorporated Statistician, 5(1):1, March 1954. CO-DEN ???? ISSN 1466-9404. URL http:

                                        //www.jstor.org/stable/2986650.

Anonymous:1954:FMa

[126] Anonymous. Front matter. The Incorporated Statistician, 5(1):2–53, March 1954. CODEN ???? ISSN 1466-9404.

URL http://www.jstor.org/stable/
2986649.

Saunders:1954:NIS

[127] C. T. Saunders. National income statistics are they accurate or useful? The Incorporated Statistician, 5(1):3–15, March 1954. CODEN ???? ISSN

1466-9404. URL http://www.jstor.
org/stable/2986651.

Downham:1954:SCS

[128] J. S. Downham. Social class in sample surveys. The Incorporated Statistician, 5(1):17–38, March 1954. CODEN ???? ISSN 1466-9404. URL http://

www.jstor.org/stable/2986652.

Hughes:1954:DDS

[129] C. Davenport Hughes. Difficulties and dangers in statistics in industry. The Incorporated Statistician, 5(1):39–49 + 51, March 1954. CODEN ???? ISSN

1466-9404. URL http://www.jstor.
org/stable/2986653.

G:1954:BRBa

[130] H. W. G. G. Book review: Elements of Statistics, by C. G. Lambe. The Incorporated Statistician, 5(1):51–52, March 1954. CODEN ???? ISSN 1466-9404.

URL http://www.jstor.org/stable/
2986654.

Anonymous:1954:CQC

[131] Anonymous. Conference at Queen’s College, Cambridge. The Incorporated Statistician, 5(1):54, March 1954. CO-DEN ???? ISSN 1466-9404. URL http:

//www.jstor.org/stable/2986655.

Anonymous:1954:BMa

[132] Anonymous. Back matter. The Incorporated Statistician, 5(1):55–56, March 1954. CODEN ???? ISSN 1466-9404.

                                        URL http://www.jstor.org/stable/
                                        2986656.

Anonymous:1954:Eb

[133] Anonymous. Editorial. The Incorporated Statistician, 5(2):57, August 1954. CO-DEN ???? ISSN 1466-9404. URL http:

                                        //www.jstor.org/stable/2986460.

Anonymous:1954:FMb

[134] Anonymous. Front matter. The Incorporated Statistician, 5(2):58–99, August 1954. CODEN ???? ISSN 1466-9404.

                                        URL http://www.jstor.org/stable/
                                        2986459.

George:1954:ICU

[135] C. Oswald George. Industrial censuses in the United States. The Incorporated Statistician, 5(2):59–82, August 1954. CODEN ???? ISSN 1466-9404.

                                        URL http://www.jstor.org/stable/
                                        2986461.

Anonymous:1954:FAG

[136] Anonymous. Fifth Annual General Meeting. The Incorporated Statistician, 5(2):84, August 1954. CODEN ???? ISSN 1466-9404. URL http://www.

                                        jstor.org/stable/2986462.

Norris:1954:SAM

[137] Vera Norris. Some aspects of medical statistics. The Incorporated Statistician, 5(2):85–98, August 1954. CODEN ???? ISSN 1466-9404. URL http://

                                        www.jstor.org/stable/2986463.

G:1954:BRBb

[138] G. Book review: Sampling Methods for Censuses and Surveys: Second Edition. The Incorporated Statistician, 5(2):100, August 1954. CODEN ???? ISSN

1466-9404. URL http://www.jstor.
org/stable/2986464.

Durant:1954:GPS

[139] Henry Durant. The Gallup Poll and some of its problems. The Incorporated Statistician, 5(2):101–112, August 1954. CODEN ???? ISSN 1466-9404.

URL http://www.jstor.org/stable/
2986465.

Anonymous:1954:BMb

[140] Anonymous. Back matter. The Incorporated Statistician, 5(2):??, August 1954. CODEN ???? ISSN 1466-9404.

URL http://www.jstor.org/stable/
2986466.

Anonymous:1954:Ec

[141] Anonymous. Editorial. The Incorporated Statistician, 5(3):113, November 1954. CODEN ???? ISSN 1466-9404.

URL http://www.jstor.org/stable/
2986644.

Anonymous:1954:FMc

[142] Anonymous. Front matter. The Incorporated Statistician, 5(3):114–160, November 1954. CODEN ???? ISSN 1466-

9404. URL http://www.jstor.org/
stable/2986643.

Geary:1954:CRS

[143] R. C. Geary. The contiguity ratio and statistical mapping. The Incorporated Statistician, 5(3):115–127 + 129–146, November 1954. CODEN ???? ISSN

                                        1466-9404. URL http://www.jstor.
                                        org/stable/2986645.

Tippett:1954:SRM

[144] L. H. C. Tippett. Statistics in research and management in the cotton industry. The Incorporated Statistician, 5 (3):147–159, November 1954. CODEN ???? ISSN 1466-9404. URL http://

                                        www.jstor.org/stable/2986646.

Anonymous:1954:JEP

[145] Anonymous. July, 1954 examinations pass list. The Incorporated Statistician, 5(3):161–162, November 1954. CODEN ???? ISSN 1466-9404. URL http://

                                        www.jstor.org/stable/2986647.

Anonymous:1954:BMc

[146] Anonymous. Back matter. The Incorporated Statistician, 5(3):163–164, November 1954. CODEN ???? ISSN 1466-

                                        9404. URL http://www.jstor.org/
                                        stable/2986648.

Anonymous:1955:Ea

[147] Anonymous. Editorial. The Incorporated Statistician, 5(4):163, April 1955. CO-DEN ???? ISSN 1466-9404. URL http:

                                        //www.jstor.org/stable/2986484.

Anonymous:1955:FMa

[148] Anonymous. Front matter. The Incorporated Statistician, 5(4):164–196, April 1955. CODEN ???? ISSN 1466-9404.

                                        URL http://www.jstor.org/stable/
                                        2986483.

Stone:1955:SEP

[149] Richard Stone, J. Aitchison, and J. A. C. Brown. Some estimation problems in demand analysis. The Incorporated Statistician, 5(4):165–177, April 1955. CO-DEN ???? ISSN 1466-9404. URL http:

//www.jstor.org/stable/2986485.

Morrell:1955:LT

[150] A. J. H. Morrell. Labour turnover. The Incorporated Statistician, 5(4):179–189, April 1955. CODEN ???? ISSN 1466-

9404. URL http://www.jstor.org/
stable/2986486.

Robson:1955:RCS

[151] R. Robson. Raw cotton statistics. The Incorporated Statistician, 5(4):190–195, April 1955. CODEN ???? ISSN 1466-

9404. URL http://www.jstor.org/
stable/2986487.

Read:1955:SMC

[152] D. R. Read. Statistical methods in chemical process development. The Incorporated Statistician, 5(4):197–207, April 1955. CODEN ???? ISSN 1466-

9404. URL http://www.jstor.org/
stable/2986488.

M:1955:BRBa

[153] J. I. M. Book review: Theory of Games and Statistical Decisions, by D. Blackwell; M. A. Girshick. The Incorporated Statistician, 5(4):208, April 1955. CO-DEN ???? ISSN 1466-9404. URL http:

//www.jstor.org/stable/2986489.

M:1955:BRBb

[154] J. I. M. Book review: The Design of Social Research, by Russell L. Ackoff. The Incorporated Statistician, 5(4):208– 209, April 1955. CODEN ???? ISSN

1466-9404. URL http://www.jstor.
org/stable/2986490.

M:1955:BRBc

[155] A. J. H. M. Book review: Introduction to Mathematical Statistics. Second Edition, by P. G. Hoel. The Incorporated Statistician, 5(4):209–210, April 1955. CODEN ???? ISSN 1466-9404. URL http://

                                        www.jstor.org/stable/2986491.

Anonymous:1955:AISa

[156] Anonymous. The Association of Incorporated Statisticians. The Incorporated Statistician, 5(4):211, April 1955. CO-DEN ???? ISSN 1466-9404. URL http:

                                        //www.jstor.org/stable/2986492.

Anonymous:1955:BMa

[157] Anonymous. Back matter. The Incorporated Statistician, 5(4):212, April 1955. CODEN ???? ISSN 1466-9404.

                                        URL http://www.jstor.org/stable/
                                        2986493.

Anonymous:1955:Eb

[158] Anonymous. Editorial. The Incorporated Statistician, 6(1):1–2, July 1955. CO-DEN ???? ISSN 1466-9404. URL http:

                                        //www.jstor.org/stable/2986367.

Rudoe:1955:RTS

[159] W. Rudoe. Retail trade statistics. The Incorporated Statistician, 6(1):3–16, July 1955. CODEN ???? ISSN 1466-

                                        9404. URL http://www.jstor.org/
                                        stable/2986368.

Anonymous:1955:FMb

[160] Anonymous. Front matter. The Incorporated Statistician, 6(1):17–44, July 1955. CODEN ???? ISSN 1466-9404.

                                        URL http://www.jstor.org/stable/
                                        2986366.

Palca:1955:SSP

[161] H. Palca. Some sampling problems in agriculture. The Incorporated Statistician, 6(1):18–33, July 1955. CODEN ???? ISSN 1466-9404. URL http://

www.jstor.org/stable/2986369.

Huppert:1955:HS

[162] Eric L. Huppert. Health statistics.

The Incorporated Statistician, 6(1):35– 43, July 1955. CODEN ???? ISSN

1466-9404. URL http://www.jstor.
org/stable/2986370.

R:1955:BRB

[163] W. R. Book review: Chance and Choice, by Cardpack and Chessboard. Volume II, by Lancelot Hogben. The Incorporated Statistician, 6(1):45–46, July 1955. CODEN ???? ISSN 1466-9404.

URL http://www.jstor.org/stable/
2986371.

M:1955:BRBd

[164] A. J. H. M. Book review: Exercises in Theoretical Statistics, by M. G. Kendall. The Incorporated Statistician, 6(1):46, July 1955. CODEN ???? ISSN 1466-

9404. URL http://www.jstor.org/
stable/2986372.

Anonymous:1955:AISb

[165] Anonymous. The Association of Incorporated Statisticians Limited. The Incorporated Statistician, 6(1):47, July 1955. CODEN ???? ISSN 1466-9404.

URL http://www.jstor.org/stable/
2986373.

Anonymous:1955:AISc

[166] Anonymous. The Association of Incorporated Statisticians Limited Conference at Balliol College, Oxford. The Incorporated Statistician, 6(1):48, July 1955. CODEN ???? ISSN 1466-9404.

URL http://www.jstor.org/stable/
2986374.

Anonymous:1955:AISd

[167] Anonymous. The Association of Incorporated Statisticians. The Incorporated Statistician, 6(1):49, July 1955. CODEN ???? ISSN 1466-9404. URL http://

                                        www.jstor.org/stable/2986375.

Anonymous:1955:BMb

[168] Anonymous. Back matter. The Incorporated Statistician, 6(1):50, July 1955. CODEN ???? ISSN 1466-9404.

                                        URL http://www.jstor.org/stable/
                                        2986376.

Anonymous:1955:Ec

[169] Anonymous. Editorial. The Incorporated Statistician, 6(2):51, October 1955. CO-DEN ???? ISSN 1466-9404. URL http:

                                        //www.jstor.org/stable/2986577.

Anonymous:1955:FMc

[170] Anonymous. Front matter. The Incorporated Statistician, 6(2):52–78, October 1955. CODEN ???? ISSN 1466-9404.

                                        URL http://www.jstor.org/stable/
                                        2986576.

Ryan:1955:SSJ

[171] John Ryan. Some statistics of the Japanese economy. The Incorporated Statistician, 6(2):53–66, October 1955. CODEN ???? ISSN 1466-9404.

                                        URL http://www.jstor.org/stable/
                                        2986578.

Williams:1955:SAD

[172] G. Prys Williams. Statistical aspects of delinquency among juvenile wageearners. The Incorporated Statistician, 6(2):67–77, October 1955. CODEN ???? ISSN 1466-9404. URL http://www.

                                        jstor.org/stable/2986579.

Hughes:1955:JD

[173] C. Davenport Hughes. Juvenile delinquency. The Incorporated Statistician, 6 (2):79–82, October 1955. CODEN ???? ISSN 1466-9404. URL http://www.

jstor.org/stable/2986580.

Hendry:1955:SPM

[174] I. F. Hendry. Statistics in paper manufacture. The Incorporated Statistician, 6 (2):83–89, October 1955. CODEN ???? ISSN 1466-9404. URL http://www.

jstor.org/stable/2986581.

Anonymous:1955:JEP

[175] Anonymous. July, 1955 examinations pass list. The Incorporated Statistician, 6(2):90–91, October 1955. CODEN ???? ISSN 1466-9404. URL http://www.

jstor.org/stable/2986582.

Huppert:1955:HSC

[176] E. L. Huppert. Health statistics: [corrections and additions]. The Incorporated Statistician, 6(2):92, October 1955. CO-DEN ???? ISSN 1466-9404. URL http:

//www.jstor.org/stable/2986583.

Anonymous:1955:AISe

[177] Anonymous. The Association of Incorporated Statisticians Limited. The Incorporated Statistician, 6(2):93, October 1955. CODEN ???? ISSN 1466-9404.

URL http://www.jstor.org/stable/
2986584.

Anonymous:1955:AISf

[178] Anonymous. The Association of Incorporated Statisticians. The Incorporated Statistician, 6(2):94, October 1955. CO-DEN ???? ISSN 1466-9404. URL http:

//www.jstor.org/stable/2986585.

Anonymous:1955:BMc

[179] Anonymous. Back matter. The Incorporated Statistician, 6(2):??, October 1955. CODEN ???? ISSN 1466-9404.

                                        URL http://www.jstor.org/stable/
                                        2986586.

Anonymous:1956:Ea

[180] Anonymous. Editorial. The Incorporated Statistician, 6(3/4):95, January 1956. CODEN ???? ISSN 1466-9404.

                                        URL http://www.jstor.org/stable/
                                        2986451.

Anonymous:1956:FMa

[181] Anonymous. Front matter. The Incorporated Statistician, 6(3/4):96–132, January 1956. CODEN ???? ISSN 1466-

                                        9404. URL http://www.jstor.org/
                                        stable/2986450.

Bray:1956:EAS

[182] F. Sewell Bray. Economic accounting: A study in conceptional relationships. The Incorporated Statistician, 6(3/4):97–123, January 1956. CODEN ???? ISSN

                                        1466-9404. URL http://www.jstor.
                                        org/stable/2986452.

Wishart:1956:ABD

[183] John Wishart. An approximation to the binomial distribution. The Incorporated Statistician, 6(3/4):124–131, January 1956. CODEN ???? ISSN 1466-

                                        9404. URL http://www.jstor.org/
                                        stable/2986453.

Cauter:1956:SAC

[184] Tom Cauter. Some aspects of classification data in market research. The Incorporated Statistician, 6(3/4):133–144, January 1956. CODEN ???? ISSN

1466-9404. URL http://www.jstor.
org/stable/2986454.

Cunliffe:1956:SUS

[185] S. V. Cunliffe. Some uses of statistics in the brewing industry. The Incorporated Statistician, 6(3/4):145–158, January 1956. CODEN ???? ISSN

1466-9404. URL http://www.jstor.
org/stable/2986455.

M:1956:BRBa

[186] J. I. M. Book review: The Analysis of Family Budgets, by S. J. Prais; H. S. Houthakker. The Incorporated Statistician, 6(3/4):159–161, January 1956. CODEN ???? ISSN 1466-9404.

URL http://www.jstor.org/stable/
2986456.

R:1956:BRBa

[187] S. R. Book review: An Introduction to Stochastic Processes with Special Reference to Methods and Applications, by M. S. Bartlett. The Incorporated Statistician, 6(3/4):161–162, January 1956. CODEN ???? ISSN 1466-

9404. URL http://www.jstor.org/
stable/2986457.

Anonymous:1956:BMa

[188] Anonymous. Back matter. The Incorporated Statistician, 6(3/4):163–165, January 1956. CODEN ???? ISSN

1466-9404. URL http://www.jstor.
org/stable/2986458.

Anonymous:1956:Eb

[189] Anonymous. Editorial. The Incorporated Statistician, 7(1):1, May 1956. CODEN ???? ISSN 1466-9404. URL http://

www.jstor.org/stable/2986495.

Anonymous:1956:FMb

[190] Anonymous. Front matter. The Incorporated Statistician, 7(1):2–41, May 1956. CODEN ???? ISSN 1466-9404.

                                        URL http://www.jstor.org/stable/
                                        2986494.

Hogben:1956:PCS

[191] Lancelot Hogben. The present crisis in statistical theory. The Incorporated Statistician, 7(1):3–21, May 1956. CO-DEN ???? ISSN 1466-9404. URL http:

                                        //www.jstor.org/stable/2986496.

Fereday:1956:SSP

[192] Joy Fereday. Statistics and the study of prehistoric races. The Incorporated Statistician, 7(1):23–40, May 1956. CO-DEN ???? ISSN 1466-9404. URL http:

                                        //www.jstor.org/stable/2986497.

Anonymous:1956:CJC

[193] Anonymous. Conference at Jesus College, Oxford. The Incorporated Statistician, 7(1):42, May 1956. CODEN ???? ISSN 1466-9404. URL http://

                                        www.jstor.org/stable/2986498.

R:1956:BRBb

[194] S. R. Book review: Methods in Numerical Analysis, by Kaj. L. Nielson. The Incorporated Statistician, 7(1):43– 44, May 1956. CODEN ???? ISSN

                                        1466-9404. URL http://www.jstor.
                                        org/stable/2986499.

J:1956:BRB

[195] J. N. R. J. Book review: Experimental Design, by W. T. Federer. The Incorporated Statistician, 7(1):44–45, May 1956. CODEN ???? ISSN 1466-9404.

                                        URL http://www.jstor.org/stable/
                                        2986500.

Anonymous:1956:AISa

[196] Anonymous. The Association of Incorporated Statisticians. The Incorporated Statistician, 7(1):46, May 1956. CODEN ???? ISSN 1466-9404. URL http://

www.jstor.org/stable/2986501.

Anonymous:1956:BMb

[197] Anonymous. Back matter. The Incorporated Statistician, 7(1):??, May 1956. CODEN ???? ISSN 1466-9404.

URL http://www.jstor.org/stable/
2986502.

Anonymous:1956:Ec

[198] Anonymous. Editorial. The Incorporated Statistician, 7(2):47, July 1956. CODEN ???? ISSN 1466-9404. URL http://

www.jstor.org/stable/2986537.

Anonymous:1956:FMc

[199] Anonymous. Front matter. The Incorporated Statistician, 7(2):48–72, July 1956. CODEN ???? ISSN 1466-9404.

URL http://www.jstor.org/stable/
2986536.

Brech:1956:BE

[200] Ronald Brech. The British economy — 1956 and 1957. The Incorporated Statistician, 7(2):49–61, July 1956. CODEN ???? ISSN 1466-9404. URL http://

www.jstor.org/stable/2986538.

Liston:1956:SBR

[201] D. J. Liston. The statistician and business research. The Incorporated Statistician, 7(2):63–71, July 1956. CODEN ???? ISSN 1466-9404. URL http://

www.jstor.org/stable/2986539.

Farlie:1956:CQ

[202] D. J. Farlie. Canteen queues. The Incorporated Statistician, 7(2):73–84, July 1956. CODEN ???? ISSN 1466-9404.

URL http://www.jstor.org/stable/
2986540.

M:1956:BRBb

[203] J. I. M. Book review: National Income Statistics; Sources, and Methods. The Incorporated Statistician, 7(2):85– 86, July 1956. CODEN ???? ISSN

1466-9404. URL http://www.jstor.
org/stable/2986541.

Anonymous:1956:EPC

[204] Anonymous. Erratum: The present crisis in statistical theory. The Incorporated Statistician, 7(2):87, July 1956. CODEN ???? ISSN 1466-9404. URL http://

www.jstor.org/stable/2986542.

Anonymous:1956:AISb

[205] Anonymous. The Association of Incorporated Statisticians. The Incorporated Statistician, 7(2):88, July 1956. CODEN ???? ISSN 1466-9404. URL http://

www.jstor.org/stable/2986543.

Anonymous:1956:BMc

[206] Anonymous. Back matter. The Incorporated Statistician, 7(2):??, July 1956. CODEN ???? ISSN 1466-9404.

URL http://www.jstor.org/stable/
2986544.

Anonymous:1956:Ed

[207] Anonymous. Editorial. The Incorporated Statistician, 7(3):89–90, December 1956. CODEN ???? ISSN 1466-9404.

URL http://www.jstor.org/stable/
2986546.

Moroney:1956:DEP

[208] M. J. Moroney. The design of experiments. Part I — the scientific background. The Incorporated Statistician, 7(3):91–99, December 1956. CODEN ???? ISSN 1466-9404. URL http://

www.jstor.org/stable/2986547.

Hunt:1956:BFA

[209] K. E. Hunt. Business forecasting as applied to agriculture. The Incorporated Statistician, 7(3):100–106, December 1956. CODEN ???? ISSN 1466-

9404. URL http://www.jstor.org/
stable/2986548.

Anonymous:1956:FMd

[210] Anonymous. Front matter. The Incorporated Statistician, 7(3):107–118, December 1956. CODEN ???? ISSN

1466-9404. URL http://www.jstor.
org/stable/2986545.

Downham:1956:MRC

[211] J. S. Downham and J. A. P. Treasure. Market research and consumer durables. The Incorporated Statistician, 7(3):108–117, December 1956. CODEN ???? ISSN 1466-9404. URL http://

www.jstor.org/stable/2986549.

I:1956:BRB

[212] A. R. I. Book review: An Introduction to British Economic Statistics, by E. Devons. The Incorporated Statistician, 7 (3):119–120, December 1956. CODEN ???? ISSN 1466-9404. URL http://

www.jstor.org/stable/2986550.

E:1956:BRB

[213] L. R. E. Book review: Readings in Market Research. The Incorporated Statistician, 7(3):120–121, December 1956. CO-DEN ???? ISSN 1466-9404. URL http:

                                        //www.jstor.org/stable/2986551.

M:1956:BRBc

[214] J. M. Book review: Industrial Censuses in the United States. The Incorporated Statistician, 7(3):121, December 1956. CODEN ???? ISSN 1466-9404.

                                        URL http://www.jstor.org/stable/
                                        2986552.

Anonymous:1956:Ee

[215] Anonymous. Education. The Incorporated Statistician, 7(3):122–124, December 1956. CODEN ???? ISSN

                                        1466-9404. URL http://www.jstor.
                                        org/stable/2986553.

Anonymous:1956:AMS

[216] Anonymous. The Air Ministry Statistics Service. The Incorporated Statistician, 7 (3):124–126, December 1956. CODEN ???? ISSN 1466-9404. URL http://

                                        www.jstor.org/stable/2986554.

Anonymous:1956:MSA

[217] Anonymous. Minutes of the Seventh Annual Meeting. The Incorporated Statistician, 7(3):126–127, December 1956. CO-DEN ???? ISSN 1466-9404. URL http:

                                        //www.jstor.org/stable/2986555.

Anonymous:1956:BMd

[218] Anonymous. Back matter. The Incorporated Statistician, 7(3):128, December 1956. CODEN ???? ISSN 1466-9404.

                                        URL http://www.jstor.org/stable/
                                        2986556.

Anonymous:1957:Ea

[219] Anonymous. Editorial. The Incorporated Statistician, 7(4):129, April 1957. CO-DEN ???? ISSN 1466-9404. URL http:

//www.jstor.org/stable/2986421.

Anonymous:1957:FMa

[220] Anonymous. Front matter. The Incorporated Statistician, 7(4):130–166, April 1957. CODEN ???? ISSN 1466-9404.

URL http://www.jstor.org/stable/
2986420.

Reece:1957:CAS

[221] J. A. Reece. Cost accounting and statistics. The Incorporated Statistician, 7 (4):131–141, April 1957. CODEN ???? ISSN 1466-9404. URL http://www.

jstor.org/stable/2986422.

Moroney:1957:DEPa

[222] M. J. Moroney. The design of experiments. Part II—The nature of measurements. The Incorporated Statistician, 7(4):143–159, April 1957. CODEN ???? ISSN 1466-9404. URL http://

www.jstor.org/stable/2986423.

Anonymous:1957:GID

[223] Anonymous. The generation of income distributions: Some illustrative models. The Incorporated Statistician, 7(4):161– 165, April 1957. CODEN ???? ISSN

1466-9404. URL http://www.jstor.
org/stable/2986424.

Anonymous:1957:DRE

[224] Anonymous. Diagrams for rapid estimation of partial correlation co-efficients. The Incorporated Statistician, 7(4):167– 169, April 1957. CODEN ???? ISSN

1466-9404. URL http://www.jstor.
org/stable/2986425.

B:1957:BRB

[225] W. R. B. Book review: Statistics — A New Approach, by W. Allen Wallis; Harry V. Roberts. The Incorporated Statistician, 7(4):170–171, April 1957. CODEN ???? ISSN 1466-9404.

URL http://www.jstor.org/stable/
2986426.

A:1957:BRB

[226] H. V. A. Book review: Statistics in Theory and Practice, by L. R. Connor; A. J. H. Morrell. The Incorporated Statistician, 7(4):171–172, April 1957. CODEN ???? ISSN 1466-9404. URL http://

www.jstor.org/stable/2986427.

Anonymous:1957:ACB

[227] Anonymous. Annual Conference at Balliol College, Oxford. The Incorporated Statistician, 7(4):173, April 1957. CO-DEN ???? ISSN 1466-9404. URL http:

//www.jstor.org/stable/2986428.

Anonymous:1957:AISa

[228] Anonymous. The Association of Incorporated Statisticians. The Incorporated Statistician, 7(4):174, April 1957. CO-DEN ???? ISSN 1466-9404. URL http:

//www.jstor.org/stable/2986429.

Anonymous:1957:BMa

[229] Anonymous. Back matter. The Incorporated Statistician, 7(4):??, April 1957. CODEN ???? ISSN 1466-9404.

URL http://www.jstor.org/stable/
2986430.

Anonymous:1957:Eb

[230] Anonymous. Editorial. The Incorporated Statistician, 8(1):1, October 1957. CO-DEN ???? ISSN 1466-9404. URL http:

//www.jstor.org/stable/2986468.

Anonymous:1957:FMb

[231] Anonymous. Front matter. The Incorporated Statistician, 8(1):2–20, October 1957. CODEN ???? ISSN 1466-9404.

URL http://www.jstor.org/stable/
2986467.

Beveridge:1957:WIP

[232] Lord Beveridge. Wages and inflation in the past. The Incorporated Statistician, 8(1):3–7, October 1957. CODEN ???? ISSN 1466-9404. URL http://

www.jstor.org/stable/2986469.

Brech:1957:AEF

[233] Ronald Brech. Assessing the efficiency of a firm. The Incorporated Statistician, 8(1):9–19, October 1957. CODEN ???? ISSN 1466-9404. URL http://

www.jstor.org/stable/2986470.

England:1957:DCS

[234] L. R. England. Department and chain store customers. The Incorporated Statistician, 8(1):21–27, October 1957. CODEN ???? ISSN 1466-9404.

URL http://www.jstor.org/stable/
2986471.

Moroney:1957:DEPb

[235] M. J. Moroney. The design of experiments. Part III. simple designs. The Incorporated Statistician, 8(1):28–37, October 1957. CODEN ???? ISSN 1466-

9404. URL http://www.jstor.org/
stable/2986472.

A:1957:CSR

[236] H. V. A. Courses in statistics at the regent street polytechnic. The Incorporated Statistician, 8(1):38–40, October 1957. CODEN ???? ISSN 1466-9404.

                                        URL http://www.jstor.org/stable/
                                        2986473.

Anonymous:1957:AGM

[237] Anonymous. Annual General Meeting. The Incorporated Statistician, 8(1):41– 43, October 1957. CODEN ???? ISSN

                                        1466-9404. URL http://www.jstor.
                                        org/stable/2986474.

Anonymous:1957:AMS

[238] Anonymous. Air Ministry Statistical Service. The Incorporated Statistician, 8(1):43, October 1957. CODEN ???? ISSN 1466-9404. URL http://www.

                                        jstor.org/stable/2986475.

Anonymous:1957:Ec

[239] Anonymous. Education. The Incorporated Statistician, 8(1):44, October 1957. CODEN ???? ISSN 1466-9404.

                                        URL http://www.jstor.org/stable/
                                        2986476.

Anonymous:1957:AC

[240] Anonymous. Annual Conference. The Incorporated Statistician, 8(1):45–46, October 1957. CODEN ???? ISSN

                                        1466-9404. URL http://www.jstor.
                                        org/stable/2986477.

H:1957:BRB

[241] N. K. H. Book review: Statistical Sampling for Auditors and Accountants, by Laurence L. Vance; John Neter. The Incorporated Statistician, 8(1):47–49, October 1957. CODEN ???? ISSN 1466-

                                        9404. URL http://www.jstor.org/
                                        stable/2986478.

G:1957:BRBa

[242] C. O. G. Book review: Annual Abstract of Statistics, No. 93, 1956, by Central Statistical Office. The Incorporated Statistician, 8(1):49–50, October 1957. CODEN ???? ISSN 1466-9404.

URL http://www.jstor.org/stable/
2986479.

G:1957:BRBb

[243] C. O. G. Book review: A Dictionary of Statistical Terms, by Maurice G. Kendall; William R. Buckland; Oliver and Boyd. The Incorporated Statistician, 8(1):50–51, October 1957. CODEN ???? ISSN 1466-9404. URL http://

www.jstor.org/stable/2986480.

Anonymous:1957:AISb

[244] Anonymous. The Association of Incorporated Statisticians. The Incorporated Statistician, 8(1):52, October 1957. CO-DEN ???? ISSN 1466-9404. URL http:

//www.jstor.org/stable/2986481.

Anonymous:1957:BMb

[245] Anonymous. Back matter. The Incorporated Statistician, 8(1):??, October 1957. CODEN ???? ISSN 1466-9404.

URL http://www.jstor.org/stable/
2986482.

Anonymous:1958:Ea

[246] Anonymous. Editorial. The Incorporated Statistician, 8(2):53, January 1958. CO-DEN ???? ISSN 1466-9404. URL http:

//www.jstor.org/stable/2986558.

Anonymous:1958:FMa

[247] Anonymous. Front matter. The Incorporated Statistician, 8(2):54–88, January 1958. CODEN ???? ISSN 1466-9404.

URL http://www.jstor.org/stable/
2986557.

Abrams:1958:TPP

[248] Mark Abrams. Technical problems in the I.P.A. Readership Survey. The Incorporated Statistician, 8(2):55–66, January 1958. CODEN ???? ISSN 1466-

                                        9404. URL http://www.jstor.org/
                                        stable/2986559.

Bessell:1958:IOA

[249] J. E. Bessell. An input-output analysis for intensive grassland dairy farms. The Incorporated Statistician, 8(2):67– 87, January 1958. CODEN ???? ISSN

                                        1466-9404. URL http://www.jstor.
                                        org/stable/2986560.

Jowett:1958:EDA

[250] G. H. Jowett. The exponential distribution and its applications. The Incorporated Statistician, 8(2):89–95, January 1958. CODEN ???? ISSN 1466-9404.

                                        URL http://www.jstor.org/stable/
                                        2986561.

Anonymous:1958:Eb

[251] Anonymous. Education. The Incorporated Statistician, 8(2):96–98, January 1958. CODEN ???? ISSN 1466-9404.

                                        URL http://www.jstor.org/stable/
                                        2986562.

B:1958:BRBa

[252] W. R. B. Book review: Statistical Methods in Research and Production, by O. L. Davies. The Incorporated Statistician, 8(2):99–100, January 1958. CO-DEN ???? ISSN 1466-9404. URL http:

                                        //www.jstor.org/stable/2986563.

Anonymous:1958:BMa

[253] Anonymous. Back matter. The Incorporated Statistician, 8(2):??, January 1958. CODEN ???? ISSN 1466-9404.

URL http://www.jstor.org/stable/
2986564.

Anonymous:1958:Ec

[254] Anonymous. Editorial. The Incorporated Statistician, 8(3):101, April 1958. CO-DEN ???? ISSN 1466-9404. URL http:

//www.jstor.org/stable/2986406.

Anonymous:1958:FMb

[255] Anonymous. Front matter. The Incorporated Statistician, 8(3):102–108, April 1958. CODEN ???? ISSN 1466-9404.

URL http://www.jstor.org/stable/
2986405.

Allen:1958:IDI

[256] R. G. D. Allen. Income distributions from inland revenue data. The Incorporated Statistician, 8(3):103–107, April 1958. CODEN ???? ISSN 1466-9404.

URL http://www.jstor.org/stable/
2986407.

Liddell:1958:SRT

[257] F. D. K. Liddell and D. J. Finch. “Syllables of recorded time”. The Incorporated Statistician, 8(3):109–120, April 1958. CODEN ???? ISSN 1466-9404.

URL http://www.jstor.org/stable/
2986408.

Brech:1958:FAT

[258] Ronald Brech. Figureless accounts: A theoretical excursion into management information. The Incorporated Statistician, 8(3):121–135, April 1958. CODEN ???? ISSN 1466-9404. URL http://

www.jstor.org/stable/2986409.

Anonymous:1958:ACJ

[259] Anonymous. Annual Conference, July, 1958. The Incorporated Statistician, 8 (3):136, April 1958. CODEN ???? ISSN

                                        1466-9404. URL http://www.jstor.
                                        org/stable/2986410.

Anonymous:1958:Ed

[260] Anonymous. Examinations. The Incorporated Statistician, 8(3):136–137, April 1958. CODEN ???? ISSN 1466-9404.

                                        URL http://www.jstor.org/stable/
                                        2986411.

Anonymous:1958:Na

[261] Anonymous. Notice. The Incorporated Statistician, 8(3):138, April 1958. CO-DEN ???? ISSN 1466-9404. URL http:

                                        //www.jstor.org/stable/2986412.

M:1958:BRBa

[262] J. I. M. Book review: Motivation Research, by H. Henry. The Incorporated Statistician, 8(3):139–140, April 1958. CODEN ???? ISSN 1466-9404.

                                        URL http://www.jstor.org/stable/
                                        2986413.

F:1958:BRB

[263] D. F. Book review: Experimental Designs, by W. G. Cochran; G. M. Cox. The Incorporated Statistician, 8(3):140– 143, April 1958. CODEN ???? ISSN

                                        1466-9404. URL http://www.jstor.
                                        org/stable/2986414.

M:1958:BRBb

[264] J. I. M. Book review: Women’s Measurements and Sizes, by Joint Clothing Council, Ltd. The Incorporated Statistician, 8(3):143–145, April 1958. CODEN ???? ISSN 1466-9404. URL http://

                                        www.jstor.org/stable/2986415.

B:1958:BRBb

[265] W. R. B. Book review: The Lognormal Distribution, by J. Aitchison; J. A.

C. Brown. The Incorporated Statistician, 8(3):145–146, April 1958. CODEN ???? ISSN 1466-9404. URL http://

www.jstor.org/stable/2986416.

A:1958:BRB

[266] H. V. A. and D. F. D. Book review:

Teach Yourself Statistics, by R. Goodman. The Incorporated Statistician, 8 (3):146–147, April 1958. CODEN ???? ISSN 1466-9404. URL http://www.

jstor.org/stable/2986417.

Anonymous:1958:AISa

[267] Anonymous. The Association of Incorporated Statisticians. The Incorporated Statistician, 8(3):148, April 1958. CO-DEN ???? ISSN 1466-9404. URL http:

//www.jstor.org/stable/2986418.

Anonymous:1958:BMb

[268] Anonymous. Back matter. The Incorporated Statistician, 8(3):??, April 1958. CODEN ???? ISSN 1466-9404.

URL http://www.jstor.org/stable/
2986419.

Anonymous:1958:Ee

[269] Anonymous. Editorial. The Incorporated Statistician, 8(4):149, July 1958. CO-DEN ???? ISSN 1466-9404. URL http:

//www.jstor.org/stable/2986378.

Anonymous:1958:FMc

[270] Anonymous. Front matter. The Incorporated Statistician, 8(4):150–176, July 1958. CODEN ???? ISSN 1466-9404.

URL http://www.jstor.org/stable/
2986377.

Friday:1958:DS

[271] Frank A. Friday. In defence of statistics.

The Incorporated Statistician, 8(4):151– 166, July 1958. CODEN ???? ISSN

                                        1466-9404. URL http://www.jstor.
                                        org/stable/2986379.

Liddell:1958:LDL

[272] F. D. K. Liddell. “Lies, Damned Lies, … ”. The Incorporated Statistician, 8 (4):167–175, July 1958. CODEN ???? ISSN 1466-9404. URL http://www.

                                        jstor.org/stable/2986380.

Moroney:1958:DEP

[273] M. J. Moroney. The design of experiments. Part IV — factorial design. The Incorporated Statistician, 8(4):177–188, July 1958. CODEN ???? ISSN 1466-

                                        9404. URL http://www.jstor.org/
                                        stable/2986381.

Anonymous:1958:AGM

[274] Anonymous. Annual General Meeting. The Incorporated Statistician, 8(4):189– 190, July 1958. CODEN ???? ISSN

                                        1466-9404. URL http://www.jstor.
                                        org/stable/2986382.

T:1958:BRBa

[275] C. G. T. Book review: Survey Methods in Social Investigation, by C. A. Moser. The Incorporated Statistician, 8(4):191 + 193, July 1958. CODEN ???? ISSN

                                        1466-9404. URL http://www.jstor.
                                        org/stable/2986383.

Anonymous:1958:BMc

[276] Anonymous. Back matter. The Incorporated Statistician, 8(4):192, July 1958. CODEN ???? ISSN 1466-9404.

                                        URL http://www.jstor.org/stable/
                                        2986386.

B:1958:BRBc

[277] W. R. B. Book review: Statistical Tables for Biological, Agricultural & Medical Research, by Ronald A. Fisher; F. Yates. The Incorporated Statistician, 8(4):193– 194, July 1958. CODEN ???? ISSN

1466-9404. URL http://www.jstor.
org/stable/2986384.

I:1958:BRB

[278] A. R. I. Book review: The Development of the Soviet Budgetary System, by R. W. Davies. The Incorporated Statistician, 8(4):194–196, July 1958. CODEN ???? ISSN 1466-9404. URL http://

www.jstor.org/stable/2986385.

Anonymous:1958:Ef

[279] Anonymous. Editorial. The Incorporated Statistician, 9(1):1, October 1958. CO-DEN ???? ISSN 1466-9404. URL http:

//www.jstor.org/stable/2986432.

Anonymous:1958:FMd

[280] Anonymous. Front matter. The Incorporated Statistician, 9(1):2, October 1958. CODEN ???? ISSN 1466-9404.

URL http://www.jstor.org/stable/
2986431.

David:1958:ASA

[281] S. T. David. An application of statistical analysis to marketing strategy. The Incorporated Statistician, 9(1):3–7, October 1958. CODEN ???? ISSN

1466-9404. URL http://www.jstor.
org/stable/2986433.

Smith:1958:TAS

[282] John S. Smith. The “G” technique of attitude scaling. The Incorporated Statistician, 9(1):9–16, October 1958. CODEN ???? ISSN 1466-9404. URL http://

www.jstor.org/stable/2986434.

Anonymous:1958:DQ

[283] Anonymous. The difficulty of quality. The Incorporated Statistician, 9(1):17– 19, October 1958. CODEN ???? ISSN

                                        1466-9404. URL http://www.jstor.
                                        org/stable/2986435.

Anonymous:1958:AC

[284] Anonymous. Annual Conference, 1958. The Incorporated Statistician, 9(1):20– 21, October 1958. CODEN ???? ISSN

                                        1466-9404. URL http://www.jstor.
                                        org/stable/2986436.

Anonymous:1958:Eg

[285] Anonymous. Education. The Incorporated Statistician, 9(1):22–26, October 1958. CODEN ???? ISSN 1466-9404.

                                        URL http://www.jstor.org/stable/
                                        2986437.

Anonymous:1958:Nb

[286] Anonymous. Notices. The Incorporated Statistician, 9(1):27–28, October 1958. CODEN ???? ISSN 1466-9404.

                                        URL http://www.jstor.org/stable/
                                        2986438.

T:1958:BRBb

[287] E. H. T. Book review: A First Course in Statistics, by R. Loveday. The Incorporated Statistician, 9(1):29–30, October 1958. CODEN ???? ISSN 1466-9404.

                                        URL http://www.jstor.org/stable/
                                        2986439.

Anonymous:1958:BMd

[288] Anonymous. Back matter. The Incorporated Statistician, 9(1):31, October 1958. CODEN ???? ISSN 1466-9404.

                                        URL http://www.jstor.org/stable/
                                        2986441.

Anonymous:1958:AISb

[289] Anonymous. The Association of Incorporated Statisticians. The Incorporated Statistician, 9(1):32, October 1958. CO-DEN ???? ISSN 1466-9404. URL http:

//www.jstor.org/stable/2986440.

Anonymous:1959:Ea

[290] Anonymous. Editorial. The Incorporated Statistician, 9(2):33, January 1959. CO-DEN ???? ISSN 1466-9404. URL http:

//www.jstor.org/stable/2986612.

Anonymous:1959:FMa

[291] Anonymous. Front matter. The Incorporated Statistician, 9(2):34–54, January 1959. CODEN ???? ISSN 1466-9404.

URL http://www.jstor.org/stable/
2986611.

Knowles:1959:RDA

[292] James W. Knowles. Recent developments in American economic statistics. The Incorporated Statistician, 9(2):35– 46, January 1959. CODEN ???? ISSN

1466-9404. URL http://www.jstor.
org/stable/2986613.

Farlie:1959:NMD

[293] D. J. G. Farlie. A note on a method for determining confidence intervals for the median. The Incorporated Statistician, 9(2):47–53, January 1959. CODEN ???? ISSN 1466-9404. URL http://

www.jstor.org/stable/2986614.

Brech:1959:SF

[294] Ronald Brech. The Statistician’s future.

The Incorporated Statistician, 9(2):55– 62, January 1959. CODEN ???? ISSN

1466-9404. URL http://www.jstor.
org/stable/2986615.

Anonymous:1959:Eb

[295] Anonymous. Education. The Incorporated Statistician, 9(2):63–65, January 1959. CODEN ???? ISSN 1466-9404.

URL http://www.jstor.org/stable/
2986616.

R:1959:BRB

[296] D. R. R. Book review: The Fundamentals of Statistical Reasoning, by M. H. Quenouille. The Incorporated Statistician, 9(2):66–67, January 1959. CODEN ???? ISSN 1466-9404. URL http://

www.jstor.org/stable/2986617.

J:1959:BRB

[297] N. L. J. Book review: An Introduction to Multivariate Statistical Analysis, by T. W. Anderson; Some Aspects of Multivariate Analysis, by S. N. Roy. The Incorporated Statistician, 9(2):67– 68, January 1959. CODEN ???? ISSN

1466-9404. URL http://www.jstor.
org/stable/2986618.

N:1959:BRB

[298] T. N. Book review: Principles of Statistical Techniques, by P. G. Moore. The Incorporated Statistician, 9(2):68– 69, January 1959. CODEN ???? ISSN

1466-9404. URL http://www.jstor.
org/stable/2986619.

B:1959:BRBa

[299] W. R. B. Book review: Applied Statistics for Engineers, by William Volk. The Incorporated Statistician, 9(2):69–71, January 1959. CODEN ???? ISSN 1466-

9404. URL http://www.jstor.org/
stable/2986620.

Anonymous:1959:AIS

[300] Anonymous. The Association of Incorporated Statisticians. The Incorporated Statistician, 9(2):72, January 1959. CO-DEN ???? ISSN 1466-9404. URL http:

//www.jstor.org/stable/2986621.

Anonymous:1959:BMa

[301] Anonymous. Back matter. The Incorporated Statistician, 9(2):??, January 1959. CODEN ???? ISSN 1466-9404.

URL http://www.jstor.org/stable/
2986622.

Anonymous:1959:Ec

[302] Anonymous. Editorial. The Incorporated Statistician, 9(3):73, April 1959. CO-DEN ???? ISSN 1466-9404. URL http:

//www.jstor.org/stable/2986566.

Anonymous:1959:FMb

[303] Anonymous. Front matter. The Incorporated Statistician, 9(3):74–96, April 1959. CODEN ???? ISSN 1466-9404.

URL http://www.jstor.org/stable/
2986565.

Farlie:1959:OCA

[304] D. J. Farlie. The operating characteristic of attribute sampling schemes. The Incorporated Statistician, 9(3):75– 84, April 1959. CODEN ???? ISSN

1466-9404. URL http://www.jstor.
org/stable/2986567.

James:1959:SV

[305] A. N. James. Sampling by variables.

The Incorporated Statistician, 9(3):85– 89, April 1959. CODEN ???? ISSN

1466-9404. URL http://www.jstor.
org/stable/2986568.

Keen:1959:TMU

[306] J. Keen and E. A. Fielding. Tendering for materials of uncertain quality. The Incorporated Statistician, 9(3):91– 95, April 1959. CODEN ???? ISSN

1466-9404. URL http://www.jstor.
org/stable/2986569.

Pearce:1959:SAM

[307] A. G. Pearce. Statistics as an aid to management in engineering workshops. The Incorporated Statistician, 9(3):97– 102, April 1959. CODEN ???? ISSN

1466-9404. URL http://www.jstor.
org/stable/2986570.

Anonymous:1959:Ed

[308] Anonymous. Education. The Incorporated Statistician, 9(3):103–106, April 1959. CODEN ???? ISSN 1466-9404.

URL http://www.jstor.org/stable/
2986571.

B:1959:BRBb

[309] W. R. B. Book review: Design and Analysis of Industrial Experiments, by O. L. Davies. The Incorporated Statistician, 9(3):107–108, April 1959. CODEN ???? ISSN 1466-9404. URL http://

www.jstor.org/stable/2986572.

I:1959:BRB

[310] A. R. I. Book review: Sampling Opinions: An Analysis of Survey Procedure, by Frederick F. Stephan; Philip J. Mc-Carthy. The Incorporated Statistician, 9 (3):108–109, April 1959. CODEN ???? ISSN 1466-9404. URL http://www.

jstor.org/stable/2986573.

G:1959:BRB

[311] T. K. G. Book review: New Developments in Industrial Control, by Martin D. Callender. The Incorporated Statistician, 9(3):109–112, April 1959. CODEN ???? ISSN 1466-9404. URL http://

www.jstor.org/stable/2986574.

Anonymous:1959:BMb

[312] Anonymous. Back matter. The Incorporated Statistician, 9(3):??, April 1959. CODEN ???? ISSN 1466-9404.

URL http://www.jstor.org/stable/
2986575.

Anonymous:1959:Ee

[313] Anonymous. Editorial. The Incorporated Statistician, 9(4):113, October 1959. CODEN ???? ISSN 1466-9404.

URL http://www.jstor.org/stable/
2988318.

Anonymous:1959:FMc

[314] Anonymous. Front matter. The Incorporated Statistician, 9(4):114–142, October 1959. CODEN ???? ISSN 1466-

9404. URL http://www.jstor.org/
stable/2988317.

Thrift:1959:NSR

[315] H. J. Thrift. The Newspaper Society Regional Readership Survey. The Incorporated Statistician, 9(4):115–137, October 1959. CODEN ???? ISSN 1466-

9404. URL http://www.jstor.org/
stable/2988319.

Ward:1959:NCP

[316] Denis H. Ward. A nomogram for calculating partial correlation coefficients. The Incorporated Statistician, 9(4):139– 141, October 1959. CODEN ???? ISSN

1466-9404. URL http://www.jstor.
org/stable/2988320.

Jowett:1959:PET

[317] G. H. Jowett. Practical experimentation as a teaching method in statistics. The Incorporated Statistician, 9(4):143– 153, October 1959. CODEN ???? ISSN

                                        1466-9404. URL http://www.jstor.
                                        org/stable/2988321.

Anonymous:1959:AGM

[318] Anonymous. Annual General Meeting. The Incorporated Statistician, 9(4):154, October 1959. CODEN ???? ISSN

                                        1466-9404. URL http://www.jstor.
                                        org/stable/2988322.

Anonymous:1959:Ef

[319] Anonymous. Education. The Incorporated Statistician, 9(4):155–158, October 1959. CODEN ???? ISSN 1466-

                                        9404. URL http://www.jstor.org/
                                        stable/2988323.

M:1959:BRB

[320] M. J. M. Book review: Experimental Designs in Industry, by V. Chew. The Incorporated Statistician, 9(4):159–160, October 1959. CODEN ???? ISSN

                                        1466-9404. URL http://www.jstor.
                                        org/stable/2988324.

Anonymous:1959:BMc

[321] Anonymous. Back matter. The Incorporated Statistician, 9(4):??, October 1959. CODEN ???? ISSN 1466-9404.

                                        URL http://www.jstor.org/stable/
                                        2988325.